# Certified Randomness implies Secure Classical Position-Verification

### v2

@@ Abstract
Liu et al. (ITCS22) initiated the study of designing a secure position verification protocol based on a specific proof of quantumness protocol and classical communication. In this paper, we study this interesting topic further and answer some of the open questions that are left in that paper. We provide a new generic compiler that can convert any single round proof of quantumness-based certified randomness protocol to a secure classical communication-based position verification scheme. Later, we extend our compiler to different kinds of multi-round proof of quantumness-based certified randomness protocols. Moreover, we instantiate our compiler with a random circuit sampling (RCS)-based certified randomness protocol proposed by Aaronson and Hung (STOC 23). RCS-based techniques are within reach of today’s NISQ devices; therefore, our design overcomes the limitation of the Liu et al. protocol that would require a fault-tolerant quantum computer to realize. Moreover, this is one of the first cryptographic applications of RCS-based techniques other than certified randomness.
@@ 1 Introduction
Secure positioning is a sub-domain of information theory where cryptographic protocols use the geographic location of an entity as the only credential. Although it was formally introduced in [CGMO09], it has been studied before in other domains such as wireless security [BC93, Bus04, CH05, SP05, SSW03, VN06, ZLFW06]. Position verification is one of the simplest and most fundamental position-based cryptographic functionalities, where the goal is to allow multiple parties to verify the geographic location of a prover.
In [CGMO09], Chandran et al. proved the impossibility of a classical position-verification scheme in the standard model. In [Ken11], Kent et al. first proposed the idea of quantum communication-based position verification under quantum tagging. Later, in [BCF+14], Buhrman et al. re-initiated the study of this topic and proved the impossibility of designing an information-theoretically secure quantum communication-based position verification protocol. The authors provide a generic strategy to attack any position verification protocol using the instantaneous non-local computation technique proposed by Vaidman [Vai03]. However, for this attack, the malicious provers need to share a doubly exponential (in the security parameter) number of EPR pairs. Later, in [BK11], the author reduced the entanglement requirement to exponential (in the security parameter) using port-based teleportation. On the other hand, in [BCF+14], Buhrman et al. showed if the adversaries do not have access to EPR pairs, then it is possible to design a secure quantum communication-based position verification scheme. Many protocols in the literature achieve security against non-entangled adversaries, but they become vulnerable to adversaries with an exponential amount of entanglement [CL15, Unr14, JKPPG22, QS15, AEFR+23]. In the standard model, the security of a quantum communication-based position verification scheme against adversaries with a polynomially bounded amount of entanglement remains open. However, in the random oracle model, Unruh [Unr14] proves the security of a quantum communication-based position verification protocol against adversaries with unbounded shared entanglement.
In [LLQ22], Liu et al. initiate the study of designing position-verification protocols that are based on proof of quantumness and classical communication. The authors show that the proofs of quantumness is necessary to design a secure classical verifier position-verification protocol. Moreover, the authors prove that if the prover has access to a fault-tolerant quantum computer, then under the LWE hardness assumption, one can design a secure classical communication-based position verification protocol, hence beating the impossibility result proposed by Chandran et al. [CGMO09].
The classical impossibility result and the results of Liu et al. make clear that the existence of proofs of quantum are a necessary condition, and that a specific proof of quantumness can act as a sufficient condition. However, it is unclear whether proofs of quantumness in general are a sufficient condition for designing a classical verifier position verification scheme (CVPV). In this paper, we answer the following question.
What fundamental properties suffice for designing a secure position verification scheme with classical verifiers?
In particular, we show a general class of proofs of quantumness which suffice to enable secure classical position verification. Moreover, the protocol proposed in [LLQ22] requires, along with the hardness of the LWE assumption, a fault-tolerant quantum computer (with quantum memory) and the LWE-hardness assumption. As such, such a protocol is beyond the reach of the current-day available noisy intermediate scale quantum (NISQ) machines. In this paper, we further pose the following question, and answer it in the affirmative.
@@
Is it possible to design a classical position verification protocol that can be implemented using a NISQ device and depends on some other computational hardness assumption?
@@ 1.1 Our Results
In this paper, we make progress toward the first question by proving that in the quantum random oracle model, a single-round protocol that achieves certified randomness with classical verifiers is sufficient for the secure position verification scheme with classical verifiers.
@@
In the literature, most of the existing certified randomness protocols are multi-round and it is not immediately clear whether we can draw a conclusion similar to Theorem 1 for the multi-round protocols. Moreover, there are many ways to extend our single-round compiler to a multi-round one, each with a different set of advantages and drawbacks. Our first multi-round compiler is a natural generalization of the single-round compiler, where we construct the position verification protocol just by composing the single-round protocols in sequence. However, this compiler works only for the multi-round protocols that achieve a stronger notion of certified randomness, namely those with the so called Sequential Decomposition property.
@@
A multi-round PoQ protocol is said to have the sequential decomposition property if the following holds: For an unbounded guesser trying to guess the prover’s answer each round and is allowed to communicate with the prover after the answer and the guess are sent to the verifiers, the guesser still fails at least one round with high probability.
Remark 1.
The above definition roughly captures the property that the security of the multi-round protocol reduces to the security of multiple iterations of the smaller single-round protocol. This should not be confused with the requirement that the protocol achieve security under common notions of sequential composition, which is an altogether separate and stronger property than the sequential decomposition property defined above.
For such multi-round certified randomness protocols, we get the following result.
@@
Assume there exists a multi-round certified randomness protocol, with a classical verifier and a computationally-bounded quantum prover, that satisfies the sequential decomposition property, then there exists a position verification protocol with classical verifiers secure in the quantum random oracle model.
Note that most of the existing multi-round certified randomness protocols with a classical verifier and computationally-bounded quantum provers are either based on the post-quantum secure trapdoor claw-free functions with an adaptive hardcore bit property or based on random circuit sampling [BCM+21a, AH23]. It is not clear whether these existing protocols would satisfy the sequential decomposition property that we need for our multi-round compiler. Although these protocols provide a lower bound on the smooth min-entropy conditioned on some side information, our setting allows communication with an guesser that is allowed to receive a copy of the classical outputs, resulting in zero entropy.
Fortunately, to be helpful in assisting the guesses, the side information for the $i$th round must be present in the $i$th round, not before or after. Although the conditional entropy cannot be lower bounded due to communication, it should not affect the guessing probability. Here, we show that protocols in Ref [BCM+21a, AH23] indeed satisfy the sequential decomposition property by proving the following theorem.
@@
If a multi-round certified randomness protocol, with a classical verifier and a computationally bounded quantum prover, has non-zero single-round von Neumann entropy “on average” when it does not abort (e.g. in [AH23, BCM+21a]), then that certified randomness protocol also satisfy the sequential decomposition property.
We refer to our second multi-round compiler as a rapid-fire compiler. Here, the verifiers send all the challenges that are related to the multi-round certified randomness protocol sequentially with a small predetermined time gap. The difference with the previous approach is that here, the verifiers do not wait for the answers to arrive from the prover before sending the next challenge. This protocol achieves the desired security in the QROM from the multi-round certified randomness protocols without the sequential decomposition property like in [BCM+21a, AH23]. We get the following result in this direction.
@@
Finally, we combine both previous approaches and provide a new compiler called Sequential Rapid-Fire Compiler that shows prospects in overcoming the limitations of the previous multi-round compilers.
@@
Later, in Appendix B, we instantiate our compilers with a random circuit sampling (RCS)–based multi-round certified randomness protocol proposed by Aaronson and Hung [AH23]. This, along with the recent proposal of [KT24] to use RCS for quantum cryptography, is one of the first cryptographic applications of RCS other than certified randomness. Related recent work [MSY24] explores the minimum complexity assumption needed for protocols such as RCS [^1] That is to say, proof of quantumness protocols with inefficient classical verification. and characterizes it as classically-secure one-way puzzles. Our construction also achieves advantages over the existing protocol in [LLQ22]. For example, compared to the protocol proposed in [LLQ22], our RCS-based protocol doesn’t need any fault-tolerant quantum computers to implement. It is within the reach of NISQ devices and, therefore, can be implemented with near-term quantum computers. Moreover, unlike the protocol in [LLQ22], our instantiation doesn’t require any quantum memory, which makes our protocol more implementation-friendly. Finally, our construction in this paper also achieves a secure position-verification scheme under the hardness assumption of random circuit sampling, which is entirely different from the LWE-hardness assumption.

### v3

# On the Equivalence between Classical Position Verification and Certified Randomness

@@ Abstract
Gate-based quantum computers hold enormous potential to accelerate classically intractable computational tasks. Random circuit sampling (RCS) is the only known task that has been able to be experimentally demonstrated using current-day NISQ devices. However, for a long time, it remained challenging to demonstrate the quantum utility of RCS on practical problems. Recently, leveraging RCS, an interactive protocol generating certified randomness was demonstrated using a trapped ion quantum computer [L+25], advancing the practical utility of near-term gate-based quantum computers. In this work, we establish a strong connection between certified randomness and another quantum computation classical communication primitive, classically verifiable position verification (CVPV), which circumvents the practical challenges that may arise from long-distance quantum communications. We provide a new generic compiler that can convert any single-round proof of quantumness based certified randomness protocol into a secure classical communication-based position verification scheme. Later, we extend our compiler to different types of multi-round protocols. Notably, our compiler can be applied to any multi-round certified randomness protocol that can be analyzed using the entropy accumulation theorem [DFR20], making its applicability very general. Moreover, we show that CVPV is equivalent to a relaxed variant of certified randomness that we define. We instantiate each of our compilers using existing certified randomness protocols. In particular, building on the work of Aaronson and Hung [AH23], we give a NISQ-friendly instantiation based on RCS, which was experimentally demonstrated by Liu et al. [L+25]. Hence, we show that CVPV is another application within reach of NISQ devices.
@@ 1 Introduction
Significant effort and investment have been devoted to the study and development of quantum computers due to their potential to accelerate a range of computational tasks that are classically intractable. Researchers have explored a myriad of applications, assessing the potential advantages [Sho94, HHL09, S+24, LAT21, BACS07]. Several problems [Sho94, JSW+24, BBK+23] have been theoretically shown to achieve exponential quantum speedup over the best-known classical algorithms. More recently, researchers have made progress in showing the utility of quantum computational advantage, either by trying to solve problems of practical relevance [K+23] or demonstrating prototypes of quantum error correction [B+24, P+24, AC24] and scalable architectures [AR+25]. Furthermore, there exist cryptographic problems that are impossible to solve using classical methods alone. Two such primitives are certified randomness and position verification [BCF+14].
The goal of certified randomness is to classically validate the output of an inherently random quantum process. It has potential applications to many areas including cryptography, blockchain, and finance [ACC+25]. Traditionally, constructions for certified randomness have relied on the violation of Bell inequalities [AM16, CK11, P+10, S+15]. However, these protocols require distant parties to share entanglement via quantum communications, which is susceptible to channel loss and noise. To circumvent this issue, a new research direction, termed quantum computing and classical communication (QCCC), aims to replace quantum communication with quantum computers that use classical communication. Within the QCCC framework, several constructions for certified randomness have been developed [AH23, BCM+21, MVV22, YZ24].
In the near future, it is likely that most quantum computers will be available as a cloud service. Consider a client who wishes to outsource a computational task using an external quantum cloud provider. Suppose that the computational task is to be performed on sensitive data subject to strict regional regulations, whereby the data is forbidden to leave a certain region. A notable example is the General Data Protection Regulation [EC] which stipulates that personal data of the European Union cannot be stored outside that region. In such a scenario, the client would want to ensure that the cloud provider operates in the region which it claims to be operate in, lest the client should be liable in case of a cyberattack.
@@
Position-based cryptography offers a solution to this problem by using geographic location as a credential to offer enhanced security in multiparty protocols. This concept, introduced in [CGMO09], has been studied in multiple settings, such as wireless and quantum communications [BC93, Bus04, CH05, SP05, SSW03, VN06, ZLFW06]. Position verification is the most fundamental position-based cryptographic functionality, where the goal is to allow multiple parties to verify the geographic location of a prover.
Position verification using only classical resources is impossible via a generic attack [CGMO09] which targets two weaknesses: (1) clonability of messages and (2) the ability to pull out randomness from a classical algorithm. The latter refers to the fact that without loss of generality, any randomized classical algorithm first generates a uniformly random seed and then performs a deterministic computation on its input combined with the random seed. A typical quantum position verification would circumvent this attack by avoiding the first weakness using quantum communication, thanks to the no-cloning theorem. However, quantum communication creates unique challenges for position verification [QS15]. As an alternative to quantum communication, Liu et al. [LLQ22] showed how to circumvent the impossibility result using computational assumptions, namely proof of quantumness (PoQ), to avoid the second weakness. This initiated the study of position verification in the QCCC framework, hence classically verifiable position verification (CVPV). In the language of the cloud scenario above, CVPV has the benefit over regular quantum position verification that the client need not possess any quantum resources.
In the QCCC framework, most of the existing certified randomness protocols [BCM+21, YZ24], as well as the CVPV construction of [LLQ22] would require a fault-tolerant quantum computer, therefore those constructions are beyond the reach of noisy intermediate-scale quantum (NISQ) devices.
Early works to demonstrate quantum computational advantage with NISQ devices relied on sampling problems, such as random circuit sampling (RCS) [A+19, W+21, Z+22, M+24, DHL+24].
Aaronson and Hung [AH23] proposed a certified randomness protocol with inefficient verification based on RCS. This is both the only known application of RCS to a useful task, and the only certified randomness protocol that is within the reach of NISQ devices.
In this spirit, a recent work [L+25] experimentally demonstrated a practical certified randomness protocol leveraging the result of Aaronson and Hung [AH23]. Through an interactive protocol, a trapped ion quantum computer was used to generate certified randomness. In this protocol, a classical verifier sends randomly generated quantum circuits to a prover with a quantum computer. The prover performs RCS on these circuits and sends the measured bit-strings back to the verifier. If the bit-strings pass a certain statistical test, they are guaranteed to have some entropy, yielding certified randomness. Currently, this approach stands out from the other certified randomness protocols as it does not require fault-tolerant quantum devices.
Position verification does not fall short of certified randomness with regards to benefiting from the QCCC framework and NISQ-friendly construction. The protocol proposed in [LLQ22] relies on the hardness of learning with errors (LWE) assumption and requires a fault-tolerant quantum computer out of reach for NISQ devices. In this paper, we pose the following question and answer it in the affirmative:
@@
Q1: Is it possible to design a classical position verification protocol that can be implemented using a NISQ device?
To achieve this, we leverage the NISQ-friendly task of RCS. We certify the position of the prover by making it generate random circuit samples at the prescribed time and location. This is ascertained by designing the protocol such that the information required to generate the samples, i.e. the circuit description, simultaneously reaches the prover from multiple directions. The prover is then asked to send identical samples of high statistical score to all directions. Our construction paves the way for another experimental demonstration of quantum advantage following [L+25].
In this paper, we answer the following question:
Q2: What is the fundamental property needed for designing a secure position verification scheme with classical verifiers?
We show that a general class of PoQ protocols with the certified randomness property can be used to construct secure classical position verification. As a corollary, we show that CVPV exists in the random oracle model assuming one of the following: (1) classical hardness of random circuit sampling (with inefficient verifiers), (2) quantum hardness of learning with errors (LWE) problem, (3) or the Aaronson-Ambainis Conjecture [AA14]. Furthermore, in the other direction, we show that any CVPV protocol can be compiled into a certified randomness protocol.
Therefore, we show that certified randomness is necessary and sufficient for constructing a position verification protocol within the QCCC framework. This connection enables the development of new classes of position verification protocols derived from certified randomness protocols, and vice versa.
@@ 1.1 Our Results
We give several compilers to achieve CVPV using certified randomness. We show that certified randomness is also necessary for CVPV, hence a fundamental property required to construct it. Then, we show how to instantiate our compiler using a NISQ-friendly RCS-based construction, as well as theoretically significant constructions based on error-correcting codes and LWE.
1.2 Compilers
To address the second question (Q2), we prove that a single-round protocol that achieves certified randomness with classical verifiers is sufficient for the secure position verification scheme with classical verifiers in the quantum random oracle model.
@@ 1.2 Compilers
We first consider any single-round PoQ-based certified randomness protocol $\mathscr{P}=(P,V)$, where $P$ is the prover and $V$ is the verifier. It is modeled as a two-message interactive protocol. In the first message, the verifier sends a random challenge $\mathsf{ch}$ to the prover. In the second message, it gets back $\mathsf{ans}$ as a response. Later, the verifier performs a verification process $\mathsf{Ver}$ to accept or reject the response.
Figure 1: (a) A spacetime diagram describing the honest interaction between two verifiers at locations $X=0$ and $X=2$ and an honest prover located at the claimed position (for simplicity, located at $X=1$ in the diagram). The verifiers pre-share random values $x,y,k,r$ prior to the start of the protocol, and set $s=ch\oplus G_{k}(x\oplus y)$ for some family of hash functions $\{G_{k}\}$ and $ch=\mathsf{Gen}(1^{n};r)$ for $\mathsf{Gen}$ the verifier functionality of some certified randomness verifier and where $accept(\cdot)$ determines a boolean function that determines whether the verifiers accept the prover’s claim. In all spacetime figures we provide, we implicitly require that the proper timing constraints are observed in order for the verifiers to accept. (b) A spacetime diagram describing the honest interaction between two verifiers at locations $X=0$ and $X=2$ and two dishonest prover located at arbitrary positions between $V_{1}$ and $V_{2}$ (for simplicity, located at $X=0.5$ and $X=1.5$ in the diagram). The provers are allowed to communicate during the protocol but are non-signaling. Notation and behavior is otherwise analogous to (a). In particular, verifiers expect to receive messages at times determined by honest prover’s position.
Construction: In our compiler, we consider that the two verifiers (say $V_{1},V_{2}$) in the CVPV protocol have access to a family of cryptographic hash functions $\left\{G_{k}\right\}_{k\in\{0,1\}^{\lambda}}:\{0,1\}^{m}\to\{0,1\}^{n}$. In addition, they share a random hash key $k$ and two random inputs $x,y\leftarrow\{0,1\}^{m}$. At time $t=0$, $V_{1}$ sends $s:=G_{k}(x\oplus y)\oplus\mathsf{ch}$, $x,k$ to the prover $P$ at a claimed location (say $X=1$ in Figure 1) and $V_{2}$ sends $y$ to $P$. Upon receiving $x,s,k,y$ from the verifiers, the prover can recover $\mathsf{ch}$ and send back the response $\mathsf{ans}$ to the verifiers. The verifiers verify the position of the prover if the following conditions are satisfied.
  1. 1.
    Timing Requirement: Both verifiers receive at time $t=2$.
  2. 2.
    Consistency Requirement: The received answers are identical.
  3. 3.
    Certified Randomness Verification: The verification process $\mathsf{Ver}$ corresponding to the certified randomness protocol accepts the response.
Note that [Unr14] also uses the idea to hide a secret by hashing the XOR of two messages that arrive from opposite directions. The difference is that [Unr14] hides the measurement basis in a QPV protocol while sending the quantum state in the clear, whereas we hide the certified randomness challenge which is the entire input of the prover in our CVPV protocol.
In the literature, most of the existing certified randomness protocols are multi-round and it is not immediately clear whether we can draw a conclusion similar to Theorem 1 for the multi-round protocols. Below, we discuss several multi-round protocols, and we leave the detailed descriptions of these protocols in the Supplementary Material.
Our first multi-round compiler is a natural generalization of the single-round compiler. For each round of the original certified randomness protocol, we use the interactive construction just like in the single-round compiler. However, this compiler works only for the multi-round protocols that achieve a stronger notion of certified randomness, namely those with the so called Sequential Decomposition property.
@@
A multi-round PoQ protocol is said to have the sequential decomposition property if the following holds: Consider an unbounded guesser trying to guess the prover’s answer. In each round, the guesser is only allowed to communicate with the prover after the answer and the guess are sent to the verifiers. The guesser must fail at least one round with high probability.
@@
Suppose there exists a multi-round certified randomness protocol, with a classical verifier and a computationally-bounded quantum prover, that satisfies the sequential decomposition property. Then, there exists a position verification protocol with classical verifiers secure in the quantum random oracle model.
It is not clear whether these existing protocols like [BCM+21, AH23] would satisfy the sequential decomposition property that we need for our multi-round compiler. Although these protocols provide a lower bound on the smooth min-entropy conditioned on some side information, our setting allows communication with a guesser that is allowed to receive a copy of the classical outputs, resulting in zero entropy.
The key enabling insight of this work is as follows: to be helpful in assisting the guesses, the side information for the $i$th round must be present in the $i$th round, not before or after. Although the conditional entropy cannot be lower bounded due to communication, it should not affect the guessing probability. Here, we show that protocols in [BCM+21, AH23] indeed satisfy the sequential decomposition property by proving the following theorem.
@@
If a multi-round certified randomness protocol, with a classical verifier and a computationally bounded quantum prover, has non-zero single-round von Neumann entropy “on average” when it does not abort (e.g. in [AH23, BCM+21]), then that certified randomness protocol also satisfies the sequential decomposition property.
In other words, any PoQ certified randomness protocol that can use the entropy accumulation theorem (EAT) [DFR20] for the multi-round security proof satisfies this property. EAT is a powerful framework for reducing multi-round entropy bounds to single-round entropy bounds even for an adaptive adversary in the device independent setting, and has significantly reduced the complexity of theoretical analysis of many cryptographic protocols [AH23, BCM+21, AFDF+18, PAB+20]. This definition is highly permissive and allows potential future certified randomness protocols to be compiled into a position verification protocol.
In addition to the aforementioned sequential compiler, we consider a rapid-fire compiler. Here, the verifiers send all the challenges that are related to the multi-round certified randomness protocol sequentially with a small predetermined time gap. The difference with the previous approach is that here, the verifiers do not wait for the answers to arrive from the prover before sending the next challenge. This protocol achieves the desired security in the QROM from the multi-round certified randomness protocols without the sequential decomposition property like in [BCM+21, AH23]. We get the following result in this direction.
@@
Finally, we combine both previous approaches and provide a new compiler called the Sequential Rapid-Fire Compiler that shows prospects in overcoming the limitations of the previous multi-round compilers.
@@
1.3 Instantiations
Later, in Section 8.1, we instantiate our compilers with a random circuit sampling (RCS)–based multi-round certified randomness protocol proposed by Aaronson and Hung [AH23]. Our construction achieves two advantages over the existing protocol in [LLQ22]. First, it is NISQ-friendly, thereby answering the first question (Q1) in the affirmative. Second, it achieves CVPV under the hardness assumption of RCS, which is does not imply the LWE-hardness assumption used by [LLQ22].
Our single-round compiler can also be instantiated using the certified randomness protocol of Yamakawa and Zhandry [YZ24], which shows that there exists a CVPV scheme secure in the random oracle model assuming the Aaronson-Ambainis Conjecture [AA14]; the details can be found in Section 8.2.
As another example, our compiler can be instantiated using the work of Brakerski et al. [BCM+21], which yields the result of [LLQ22] that CVPV exists in the random oracle model assuming the hardness of LWE, albeit with a different construction.
1.4 Prior Work
In [CGMO09], Chandran et al. proved the impossibility of a classical position verification scheme in the standard model. In [Ken11], Kent et al. first proposed the idea of quantum communication-based position verification under quantum tagging. Later, in [BCF+14], Buhrman et al. re-initiated the study of this topic and proved the impossibility of designing an information-theoretically secure quantum communication-based position verification protocol. The authors provide a generic strategy to attack any position verification protocol using the instantaneous non-local computation technique proposed by Vaidman [Vai03]. However, for this attack, the malicious provers need to share a doubly exponential (in the security parameter) number of EPR pairs. Later, in [BK11], the author reduced the entanglement requirement to exponential (in the security parameter) using port-based teleportation. On the other hand, in [BCF+14], Buhrman et al. showed if the adversaries do not have access to EPR pairs, then it is possible to design a secure quantum communication-based position verification scheme. Many protocols in the literature achieve security against non-entangled adversaries, but they become vulnerable to adversaries with an exponential amount of entanglement [CL15, Unr14, JKPPG22, QS15, AEFR+23]. In the standard model, the security of a quantum communication-based position verification scheme against adversaries with a polynomially bounded amount of entanglement remains open. However, in the random oracle model, Unruh [Unr14] proves the security of a quantum communication-based position verification protocol against adversaries with unbounded shared entanglement. Our compilers use share techniques of query-extraction with Unruh [Unr14]. While [Unr14] makes a reduction to a monogamy-of-entanglement game in the quantum-communication setting, we make a reduction to certified randomness in the classical-communication setting.
In [LLQ22], Liu et al. initiate the study of designing position verification protocols that are based on proof of quantumness and classical communication. The authors show that the proofs of quantumness is necessary to design a secure classical verifier position verification protocol. Moreover, the authors prove that if the prover has access to a fault-tolerant quantum computer, then under the LWE hardness assumption, one can design a secure classical communication-based position verification protocol, hence beating the impossibility result proposed by Chandran et al. [CGMO09].
1.5 Discussion
Unlike position verification based on quantum communication, our RCS-based protocol uses classical communication and can be much more resilient against environmental noise encountered during communication. However, the use of a quantum computer introduces new opportunities for noise to degrade the quality of the protocol. In the RCS-based instantiation, this effect manifests as a reduction in the achievable certified randomness XEB score, which reduces the entropy guarantee.
We also highlight the remaining problems related to classical verifier position verification. In this paper, we prove the soundness of the compilers in quantum random oracle model. However, it is crucial to have the security analysis in the plain model for a practical implementation of this protocol. Additionally, this paper requires proof of quantumness-based certified randomness to construct the CVPV protocol. However, it is not clear whether this is the most minimal assumption. We leave the minimum sufficient assumption required to construct a CVPV protocol as an open problem.
Practical implementation of the RCS-based protocol also faces additional challenges. Thus far, the security model considered in the experimental demonstration of RCS-based certified randomness assumes a family of finite-size adversaries using existing state-of-the-art techniques instead of the more general asymptotic adversaries considered in [AH23]. This restriction was a result of the challenging conditions imposed on the experiment due to a combination of limiting factors, including the achievable quantum computing fidelity, latency, and the limited classical computational budget of the verifier to complete the exponentially expensive verification task. Additional work remains to close the gap between experimental feasibility and security under more general adversary models.
Further, there are additional challenges associated with position verification. First, one needs to establish line-of-sight links between the verifiers and the prover, or at least whatever routing cannot substantially increase the total communication time. Second, the timing requirement is very stringent, which is especially difficult for current quantum computing devices. If we assume that an adversary can execute the quantum circuit instantaneously, the positional uncertainty is at least the speed of light times the time it takes for an honest quantum computer to execute the quantum circuit. Further experimental progress on latency reduction is needed to make position verification practical.